$SPAD/src/input as-eg4.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)co ecfact.as
n1:=nextPrime 1000000
n2:=nextPrime n1
n:=n1*n2
LenstraEllipticMethod n

———-

References^{[1] nothing}

4